{5x + 3y= 7
{x - 6y = 8

Answer by RAY100(1637)

(1) 5x +3y = 7
.
(2) x-6y = 8
.
let (2a) = (2) * (-5) = -5x +30y = -40
.
add (1) + (2) = 0x =33y =-33,,,,or y=-1
.
subst in (2) x-6(-1) = 8,,, x+6=8,,,,x=2
.
check
(1) 10 -3 =7,,,,ok
(2) 2-(-6) =8,,,,ok